# Vec-LUT: Vector Table Lookup for Parallel Ultra-Low-Bit LLM Inference on Edge Devices

### v1

@@ 3.1. System Overview
Figure 4 shows the overview of Vec-LUT’s mpGeMM kernel. It involves two stages: offline weight transformation and runtime LUT computation. (i) At the offline stage, Vec-LUT flexibly packs floating-point weights to low-bit integers, and then permutes them to a tile-contiguous layout. The transformed weights will be directly used as lookup indices at runtime. (ii) At the runtime stage, Vec-LUT transposes feature-contiguous activations (common in inference frameworks like llama.cpp) to token-contiguous, and then perform LUT precomputing and lookup in tiles. Specifically, the vector LUT contains precomputed results from the activation (introduced in §2.1.3), and each lookup index (transformed from weights) maps to a vector of precomputed results (i.e., $1\rightarrow N$ lookup).
@@ 23  $trit\leftarrow\lfloor i/3^{j-1}\rfloor\bmod 3$;
The core difference between vector LUT and existing scalar LUT is that Vec-LUT precomputes (line 4 to 6) and looks up (line 10 to 18) the LUT in the granularity of vectors, instead of single elements. This significantly reduces random con-contiguous memory access in scalar LUT, with the technical challenges solved by §3.3 and §3.4.
@@ 3.3. Vector LUT-Centric Tensor Layout
In most inference frameworks (e.g., llama.cpp), the activations and outputs are in a feature-first layout, which is different from Vec-LUT’s token-first layout. Therefore, to integrate into these frameworks, Vec-LUT needs extra transposition and reverse-transposition for activations and outputs, respectively. To minimize the overhead of these extra operation, we fuse them into LUT precomputing and results accumulation, achieving significant end-to-end inference acceleration (up to $4.2\times$ in prefilling and $3.2\times$ in parallel decoding in §5.3).
@@ 3.4. Cache-Aware Streamed Lookup
To reduce the LUT memory footprint (hundreds of MiBs if not tiled) while reducing random access during lookup, Vec-LUT adopts a locality-aware tiling strategy. As shown in Fig. 6, it tiles the LUT by sub-tables ($K/g$) with a tile size of $K_{tile}/g$, and by tokens ($N$) with a tile size of $N_{tile}$. The $3^{g}$ rows of sub-tables are not tiled as they will be randomly accessed during table lookup. This reduces the LUT memory footprint by up to $(K/K_{tile})(N/N_{tile})\times$, and the optimal tile size depends on hardware specifications (details in §4). In our implementation, it achieves KiB level LUT size per tile, which is sufficient to fit in typical L1 cache, achieving less than 10% cache-miss overhead in VTune (Intel2025VTune) profiling.
@@
Besides the locality issues, another source of the memory pressure is the precision mismatch during accumulation. It requires accumulating low-bit (e.g., INT8) lookup results to high-precision (e.g. INT32) outputs to avoid overflow, which reduces the accumulation throughput by $4\times$ compared to INT8 accumulation. To address this, we adopt a hierarchical accumulation design that uses INT16 to bridge INT8 and INT32: (i) LUT precomputing from INT8 (activations) to INT16 (LUT elements). (ii) Intra-block accumulation in INT16 (LUT elements and block-wise intermediates). It won’t cause overflow as long as the block size $B\leq\lfloor\frac{max(\text{INT16})}{max(\text{INT8})\times g}\rfloor$. (iii) Inter-block accumulation from INT16 (intermediates) to INT32 (outputs). With sufficiently large $B$ (e.g., $B=64$ when $g=4$), this hierarchy significantly improves the accumulation throughput.
@@ 4. Implementation
We implement and integrate Vec-LUT to llama.cpp, the most popular open-source LLM inference framework on edge devices with SOTA performance. Vec-LUT supports two custom lossless packing methods: I1 (b1.60) and I2 (b2.00). Due to the tensor layout mismatch, we implement the optional activation and output transformation steps, and fuse them with LUT precomputing and accumulation, as described in §3.3.
@@
Vec-LUT can directly calculate the optimal tile size $K_{tile}$ and $N_{tile}$ based on hardware specifications, ensuring its out-of-box efficiency without tuning.
@@
The gain of $N$ tiling comes from a reduced memory footprint while being able to utilize hardware parallelism (e.g., SIMD addition). Therefore, it is optimal to minimize $N_{tile}$ while keeping it as multiples of the maximum SIMD width in INT16 (the accumulation precision). For example, ARM NEON (128-bit) (ARM2025NEON) requires $8\mid N_{tile}$, and AVX2 (256-bit) (Intel2022AVX2) requires $16\mid N_{tile}$. In our evaluation, $N_{tile}=32$ is empirically the most performant and universal configuration.
@@ 5. Evaluation
Vec-LUT outperforms all baselines (both LUT-based and MAD-based) on tested devices, with $1.5-5.0\times$, $1.2-4.1\times$, and $1.4-3.0\times$ average speedup in the mpGeMM benchmark, end-to-end prefilling, and parallel decoding, respectively.
@@
Frame- works Packing Methods BPWs HF BitNet 3B (1bitllm2024bitnet) Llama3 8B (hf1bitllm2024llama3) Falcon3 1B (Falcon3)
@@ 5.1. Experimental Setup
We compare Vec-LUT with 3 baseline frameworks and 5 packings as listed in Table 3. (i) llama.cpp (llamacpp) is the state-of-the-art MAD-based inference framework for ternary LLMs. It has two MAD-based packings: TQ1_0 (b1.69) and TQ2_0 (b2.06). (ii) T-MAC (wei2024tmac) is the state-of-the-art LUT-based mpGeMM kernel for ternary LLMs. It uses TVM (chen2018tvm) for kernel tuning, and is integrated with llama.cpp for end-to-end inference. It has one LUT-based packing: INT_N (b2.00). (iii) bitnet.cpp (wang2025bitnet) is Microsoft’s official inference framework for ternary LLMs, built atop llama.cpp. Its LUT-based kernel is based on T-MAC. It has 3 packings supporting different devices and models: MAD-based I2_S (b2.00), and LUT-based TL1 (b2.00) and TL2 (b1.67). We fail to run bitnet.cpp (#8f75f99) on tested ARM devices, and only test it on x86 devices with supported packings.
@@
We use 3 real-world ternary LLMs to evaluate Vec-LUT. The models and packings that support them are listed in Table 3. We include both 2-bit and sub-2-bit packing for comparison as long as supported by models, frameworks, and devices. One exception is that llama.cpp falls back to Q4_0 (b4.50) quantization for HF BitNet 3B, as it only supports weights shapes divisible by 256. Although not our baseline for comparison due to the large gap in BPW ($2\times$), we still show its evaluation results.
@@
The evaluation is focused on system performance, rather than model accuracy, since our method is lossless for ternary weights. Specifically, we conduct a kernel-level mpGeMM benchmark in §5.2, and end-to-end comparisons on 3 real-world workloads in §5.3, including prefilling (§5.3.1), parallel decoding, and continuous batching that mixes prefilling and decoding (§5.3.2). We compare runs/s (higher is better) in the mpGeMM benchmark, and tokens/s (higher is better) in the end-to-end tests.
@@
Both I1 and I2 packing of Vec-LUT outperform MAD-based llama.cpp in all test cases. When comparing in similar BPWs (i.e., I1 (b1.60) vs. TQ1_0 (b1.69) and I2 (b2.00) vs. TQ2_0 (b2.06)), Vec-LUT respectively achieve $3.4-4.7\times$ and $1.7-2.9\times$ speedup in single-thread mpGeMM, and $2.9-3.7\times$ and $1.6-2.8\times$ speedup in multi-thread mpGeMM, averaged among GeMM shapes. Notably, Vec-LUT’s I1 achieves comparable or slightly higher performance than I2, while llama.cpp’s TQ1_0 is significantly slower than TQ2_0, which demonstrates the bit-scaling efficiency of Vec-LUT.
@@
Vec-LUT also outperforms scalar LUT-based T-MAC in all test cases. Specifically, Vec-LUT’s I2 (b2.00) achieves $1.8-3.2\times$ and $1.5-1.9\times$ average speedup over T-MAC’s INT_N (b2.00) in single-thread and multi-thread mpGeMM, respectively. In some cases (e.g., Intel PC and AMD laptop), T-MAC cannot outperform llama.cpp’s MAD-based TQ2_0, which proves the necessity of our vector LUT design for parallel inference.
@@
Although Vec-LUT outperforms baselines in different threads, the improvement in multi-thread is less significant than in single-thread. For example, the speedup of Vec-LUT’s I1 over llama.cpp’s TQ1_0 drops from $3.4-4.7\times$ in single-thread to $2.9-3.7\times$ in multi-thread. The main reason is that during LUT precomputing, different threads access the activation concurrently, and compete for the main memory bandwidth.
@@ 5.3. End-to-End LLM Inference
We evaluate the end-to-end inference throughput on different devices, threads, and workloads (prefilling, parallel decoding, and continuous batching). We follow llama.cpp’s practice to quantize output and token embedding weights to Q6_K and Q4_K, respectively (compilade2024tq). We denote llama.cpp’s TQ2_0 and TQ1_0 as Q4_0 explicitly when they fall back on unsupported GeMM shapes. bitnet.cpp is only evaluated on x86 devices since it fails to run on the tested ARM devices.
@@ 5.3.1. Prefilling
There are 3 MAD-based baselines: llama.cpp’s TQ1_0 (b1.69) and TQ2_0 (b2.06), and bitnet.cpp’s I2_S (2.00). Vec-LUT’s I2 (b2.00) outperforms them consistently on different devices, by $3.0-4.1\times$, $1.5-2.9\times$, and $1.6-2.3\times$ in single-thread, and $2.7-3.5\times$, $1.2-2.6\times$, and $1.2-2.2\times$ in multi-thread, averaged among different models and prompt lengths. For sub-2-bit packing specifically, Vec-LUT’s I1 (b1.60) outperforms llama.cpp’s TQ1_0 by $3.1-4.0\times$ in single-thread and $2.5-3.0\times$ in multi-thread. Despite the overall speedup (consistent with §5.2), llama.cpp’s Q4_0 on AWS Graviton 3 is an exception. The main reason is that llama.cpp’s 4-bit kernel is highly optimized on ARM devices with hardware intrinsics, outperforming both its 2-bit kernel and T-MAC. Considering the large gap in memory requirements (by $2\times$), Vec-LUT is still preferable in memory-constrained scenarios.
@@
There are 2 LUT-based baselines: T-MAC’s INT_N (b2.00) and bitnet.cpp’s TL2 (b1.67). Vec-LUT outperforms both of them, except HF BitNet 3B prefilling on the smartphone in single-thread. When comparing in similar BPWs (i.e., Vec-LUT’s I2 (b2.00) vs. T-MAC, and I1 (b1.67) vs. bitnet.cpp’s TL2), Vec-LUT respectively achieves $1.1-2.6\times$ and $1.9-2.4\times$ single-thread speedup, and $1.3-2.0\times$ and $1.7-1.8\times$ multi-thread speedup.
@@ 5.3.2. Parallel Decoding and Continuous Batching
Table 5. Vec-LUT’s latency breakdown in HF BitNet 3B prefilling. Measured with VTune (Intel2025VTune).
@@
Parallel decoding
Vec-LUT achieves similar speedup in parallel decoding as in prefilling, with up to $\times, 2.1×2.1\times$, and $3.0\times$ improvements over T-MAC, bitnet.cpp, and llama.cpp, respectively. Despite the overall improvements, we notice 2 exceptions. (i) llama.cpp’s Q4_0 still outperforms Vec-LUT for HF BitNet 3B on AWS Graviton 3, due to the same reason explained in §5.3.1. (ii) bitnet.cpp’s MAD-based I2_S (b2.00) slightly outperforms Vec-LUT on intel PC and AMD laptop with Falcon 1B. The main reason is that Falcon 1B has much smaller $M$ in its GeMM shapes (e.g., 204861442048-6144 in Falcon 1B vs. 4096144364096-14436 in Llama3 8B), which limits the benefit from table lookup. Besides, the concurrent memory access, as discussed in §5.2, also makes Vec-LUT’s multi-thread acceleration less significant.
Continuous Batching
We conduct an end-to-end serving test with continuous batching to showcase Vec-LUT’s capability to support complex parallel workloads. On an 8-core AWS Graviton 3 server that costs only \$0.50/h, Vec-LUT’s I2 can serve Falcon 1B for a 273.5 tokens/s average throughput over 32 parallel requests, including 152.6 tokens/s prefilling and 120.9 tokens/s decoding, outperforming llama.cpp’s TQ2_0 by $1.4\times$. This also demonstrates the viability of deploying ternary LLMs on affordable CPU servers, making advanced language models more accessible without requiring specialized hardware or expensive computational resources.
@@ 5.5. Ablation Study
Fig. 11 demonstrates the effectiveness of all main techniques of Vec-LUT, including vector LUT-centric tensor layout, hierarchical accumulation, streamed precomputing-lookup, topological precomputing, and locality-aware tiling. Specifically, the layout optimizations provides the initial $10.6-11.9\times$ speedup, and the streamed LUT optimizations provide another $4.1-9.0\times$ speedup.
Different input lengths
Fig. 12 shows Vec-LUT’s performance under different input lengths and $N$-tiling sizes. Unlike T-MAC’s constant throughput despite increased parallel tokens, Vec-LUT’s throughput increases linearly with the prompt length, and reaches its peak when the prompt length equals $N$-tiling size. Also, the tiling size of 32 outperforms 64. This is consistent with the analysis in §4.
@@ 6.1. Different Hardware Platforms
Following existing LUT-based works (wei2024tmac; wang2025bitnet), we demonstrate Vec-LUT’s effectiveness on the ubiquitous and LUT-friendly CPUs. It is particularly important for edge and IoT devices. Specifically, CPU has a good support for table lookup and vector addition (Vec-LUT’s main operation). Other hardware (e.g., GPU and NPU), although becoming prevalent on new generation devices, are multiplication-centric, and doesn’t suit ultra-low-bit inference. Besides, it is often impractical to use them (e.g., GPU might be reserved for rendering only), or difficult to program them (e.g., the NPU ecosystem is quite closed and tend to support fixed quantization formats). Moreover, we hope Vec-LUT to inspire future research on vector LUT-centric hardware.

### v2

@@ 3.1. System Overview
Figure 5 shows the overview of Vec-LUT’s mpGeMM kernel. It involves two stages: offline weight transformation and runtime LUT computation. (i) At the offline stage, Vec-LUT flexibly packs floating-point weights to low-bit integers, and then permutes them to a tile-contiguous layout. The transformed weights will be directly used as lookup indices at runtime. (ii) At the runtime stage, Vec-LUT transposes feature-contiguous activations (common in inference frameworks like llama.cpp) to token-contiguous, and then performs LUT precomputing and lookup in tiles. Specifically, the vector LUT contains precomputed results from the activation (introduced in §2.1.3), and each lookup index (transformed from weights) maps to a vector of precomputed results (i.e., $1\rightarrow N$ lookup).
@@ 23  $trit\leftarrow\lfloor i/3^{j-1}\rfloor\bmod 3$;
The core difference between vector LUT and existing scalar LUT is that Vec-LUT precomputes (line 4 to 6) and looks up (line 10 to 18) the LUT in the granularity of vectors, instead of single elements. This significantly reduces random non-contiguous memory access in scalar LUT, with the technical challenges solved by §3.3 and §3.4.
@@ 3.3. Vector LUT-Centric Tensor Layout
In most inference frameworks (e.g., llama.cpp), the activations and outputs are in a feature-first layout, which is different from Vec-LUT’s token-first layout. Therefore, to integrate into these frameworks, Vec-LUT needs extra transposition and reverse-transposition for activations and outputs, respectively. To minimize the overhead of these extra operations, we fuse them into LUT precomputing and results accumulation, achieving significant end-to-end inference acceleration (up to $4.2\times$ in prefilling and $3.2\times$ in parallel decoding in §5.3).
@@ 3.4. Cache-Aware Streamed Lookup
To reduce the LUT memory footprint (hundreds of MiBs if not tiled) while reducing random access during lookup, Vec-LUT adopts a locality-aware tiling strategy. As shown in Fig. 7, it tiles the LUT by sub-tables ($K/g$) with a tile size of $K_{tile}/g$, and by tokens ($N$) with a tile size of $N_{tile}$. The $3^{g}$ rows of sub-tables are not tiled as they will be randomly accessed during table lookup. This reduces the LUT memory footprint by up to $(K/K_{tile})(N/N_{tile})\times$, and the optimal tile size depends on hardware specifications (details in §4). In our implementation, it achieves KiB level LUT size per tile, which is sufficient to fit in typical L1 cache, achieving less than 10% cache-miss overhead in VTune (Intel Corporation, 2025) profiling.
@@
Besides the locality issues, another source of the memory pressure is the precision mismatch during accumulation. Following llama.cpp and the models’ official implementations on Huggingface, the activations are per-token symmetrically quantized to INT8 for input, with results accumulated in INT32 for output. This avoids overflow but reduces arithmetic throughput by $4\times$ compared to INT8. To address this, we adopt a hierarchical accumulation design that uses INT16 to bridge INT8 and INT32: (i) LUT precomputing from INT8 (activations) to INT16 (LUT elements). (ii) Intra-block accumulation in INT16 (LUT elements and block-wise intermediates). It doesn’t involve clipping or scaling, and won’t cause overflow as long as the block size $B\leq\lfloor\frac{max(\text{INT16})}{max(\text{INT8})\times g}\rfloor$. (iii) Inter-block accumulation from INT16 (intermediates) to INT32 (outputs). With sufficiently large $B$ (e.g., $B=64$ when $g=4$), this hierarchy significantly improves accumulation throughput, without introducing numerical error.
@@ 4. Implementation
We implement and integrate Vec-LUT into llama.cpp, the most popular open-source LLM inference framework on edge devices with SOTA performance. Vec-LUT supports two lossless packing methods: I1 (b1.60) and I2 (b2.00). We integrate them as two custom quantization types in llama.cpp and implement our optimizations within the mpGeMM kernel, without modifying the high-level inference flow. Due to the tensor layout mismatch, we implement the optional activation and output transformation steps, and fuse them with LUT precomputing and accumulation, as described in §3.3.
@@
Vec-LUT determines all system parameters offline, including group size $g$, accumulation block size $B$, and tile sizes $N_{tile}$ and $K_{tile}$. Specifically, Vec-LUT can directly calculate the optimal tile size $K_{tile}$ and $N_{tile}$ based on hardware specifications, ensuring its out-of-box efficiency without tuning.
@@
The gain of $N$ tiling comes from a reduced memory footprint while being able to utilize hardware parallelism (e.g., SIMD addition). Therefore, it is optimal to minimize $N_{tile}$ while keeping it as multiples of the maximum SIMD width in INT16 (the accumulation precision). For example, ARM NEON (128-bit) (Arm Limited, 2025a) requires $8\mid N_{tile}$, and AVX2 (256-bit) (Intel Corporation, 2022b) requires $16\mid N_{tile}$. In our evaluation, $N_{tile}=32$ is empirically the most performant and universal configuration.
@@ 5. Evaluation
[…]
@@
Frame- works Packing Methods BPWs HF BitNet 3B (1bitLLM, 2024) Llama3 8B (Mekkouri et al., 2024) Falcon3 1B (Falcon-LLM Team, 2024)
@@ 5.1. Experimental Setup
We compare Vec-LUT with 3 baseline frameworks  [^5] Unless otherwise noted, we use llama.cpp (#833e2b7), T-MAC (#8f29b96), and bitnet.cpp (#8f75f99) in our evaluation; llama.cpp (#5748be5), with Hexagon NPU support, is used specifically for a comparison with NPU (§5.7). and 5 packings as listed in Table 3. (i) llama.cpp (ggml-org, 2025) is the state-of-the-art MAD-based inference framework for ternary LLMs. It has two MAD-based packings: TQ1_0 (b1.69) and TQ2_0 (b2.06). (ii) T-MAC (Wei et al., 2025) is the state-of-the-art LUT-based mpGeMM kernel for ternary LLMs. It uses TVM (Chen et al., 2018) for kernel tuning, and is integrated with llama.cpp for end-to-end inference. It has one LUT-based packing: INT_N (b2.00). (iii) bitnet.cpp (Wang et al., 2025) is Microsoft’s official inference framework for ternary LLMs, built atop llama.cpp. Its LUT-based kernel is based on T-MAC. It has 3 packings supporting different devices and models: MAD-based I2_S (b2.00), and LUT-based TL1 (b2.00) and TL2 (b1.67). We fail to run bitnet.cpp (#8f75f99) on tested ARM devices, and only test it on x86 devices with supported packings.
@@
We use 3 real-world ternary LLMs to evaluate Vec-LUT. The models and packings that support them are listed in Table 3. We include both 2-bit and sub-2-bit packing for comparison as long as supported by models, frameworks, and devices. One exception is that llama.cpp falls back to Q4_0 (b4.50) quantization for HF BitNet 3B, as it only supports weight shapes divisible by 256. Although not our baseline for comparison due to the large gap in BPW ($2\times$), we still show its evaluation results.
@@
The evaluation is focused on system performance, rather than model accuracy, since our method is lossless for ternary weights. The main evaluation includes a kernel benchmark and 3 end-to-end comparisons. In §5.2, we compare operator throughput (runs/s, higher is better) of the mpGeMM kernel. In §5.3, we compare token throughput (tokens/s, higher is better) on 3 real-world workloads, including prefilling (§5.3.1), parallel decoding (§5.3.2), and continuous batching (§5.3.2) that dynamically mixes prefilling and decoding. Additionally, we compare energy efficiency (tokens/Joule, higher is better) in §5.6.
@@
[…]
@@
[…]
@@
[…]
@@ 5.3. End-to-End LLM Inference
We evaluate the end-to-end inference throughput on different devices, threads, and workloads (prefilling, parallel decoding, and continuous batching). We follow llama.cpp’s practice to quantize output and token embedding weights to Q6_K and Q4_K, respectively (compilade, 2024). We denote llama.cpp’s TQ2_0 and TQ1_0 as Q4_0 explicitly when they fall back on unsupported GeMM shapes. bitnet.cpp is only evaluated on x86 devices since it fails to run on the tested ARM devices.
@@ 5.3.1. Prefilling
[…]
@@
[…]
@@ 5.3.2. Parallel Decoding and Continuous Batching
Parallel decoding
Vec-LUT achieves similar speedup in parallel decoding as in prefilling, with up to $\times, 2.1×2.1\times$, and $3.0\times$ improvements over T-MAC, bitnet.cpp, and llama.cpp, respectively. Despite the overall improvements, we notice 2 exceptions. (i) llama.cpp’s Q4_0 still outperforms Vec-LUT for HF BitNet 3B on AWS Graviton 3, due to the same reason explained in §5.3.1. (ii) bitnet.cpp’s MAD-based I2_S (b2.00) slightly outperforms Vec-LUT on Intel PC and AMD laptop with Falcon 1B. The main reason is that Falcon 1B has much smaller $M$ in its GeMM shapes (e.g., 204861442048-6144 in Falcon 1B vs. 4096144364096-14436 in Llama3 8B), which limits the benefit from table lookup, for both vector and scalar LUT. In practice, the switching point can be determined by profiling. Besides, the concurrent memory access, as discussed in §5.2, also makes Vec-LUT’s multi-thread acceleration less significant.
Continuous Batching
We conduct an end-to-end serving test with continuous batching to showcase Vec-LUT’s capability to support complex parallel workloads with mixed prefilling and decoding in dynamic batches. On an 8-core AWS Graviton 3 server that costs only \$0.50/h, Vec-LUT’s I2 can serve Falcon 1B for a 273.5 tokens/s average throughput over 32 parallel requests, including 152.6 tokens/s prefilling and 120.9 tokens/s decoding, outperforming llama.cpp’s TQ2_0 by $1.4\times$. This also demonstrates the viability of deploying ternary LLMs on affordable CPU servers, making advanced language models more accessible without requiring specialized hardware or expensive computational resources.
@@
Table 5. Vec-LUT’s latency breakdown in HF BitNet 3B prefilling. Measured with VTune (Intel Corporation, 2025).
@@
Table 6. Energy efficiency (tokens/Joule) on AMD Ryzen 7 5800H with Llama3 8B (prefilling 256 tokens): Vec-LUT (I2) vs. llama.cpp (TQ2_0) and T-MAC (INT_N).
#Threads llama.cpp T-MAC Vec-LUT
1 0.80 1.13 1.70
4 1.32 1.40 2.54
Table 7. Throughput (tokens/s) on Snapdragon 8 Elite with Llama3 8B (prefilling 32/128 tokens): Vec-LUT (I2, CPU) vs. llama.cpp’s Hexagon backend (Q4_0, NPU).
#Tokens llama.cpp Vec-LUT Vec-LUT
(NPU) (single-thread) (dual-thread)
32 26.79 19.38 30.11
128 24.89 18.46 26.05
@@ 5.5. Ablation Study
[…]
Different input lengths and tile sizes
Fig. 13 shows Vec-LUT’s performance under different input lengths and $N$-tiling sizes. Unlike T-MAC’s constant throughput despite increased parallel tokens, Vec-LUT’s throughput increases linearly with the prompt length, surpassing T-MAC at $N=8$ ($N_{tile}=32$),and reaches its peak when the prompt length equals $N$-tiling size. Also, the tile size of 32 outperforms 64, consistent with the analysis in §4.
5.6. Energy Efficiency
[…]
5.7. Comparison with NPU
[…]
@@ 6.1. Different Hardware Platforms
Following existing LUT-based works (Wei et al., 2025; Wang et al., 2025), we demonstrate Vec-LUT’s effectiveness on the ubiquitous and LUT-friendly CPU. It is particularly important for edge and IoT devices. Specifically, CPU has a good support for table lookup and vector addition (Vec-LUT’s main operations). Other hardware (e.g., GPU and NPU), although becoming prevalent on new-generation devices, are multiplication-centric, and don’t suit ultra-low-bit inference. Besides, it is often impractical to use them (e.g., GPU might be reserved for rendering only), or difficult to program them (e.g., the NPU ecosystem is quite closed and tends to support fixed quantization formats). Moreover, we hope Vec-LUT will inspire future research on vector LUT-centric hardware.